# and non-Abelian anyons

### v1

@@ Abstract
We introduce a frequency–tunable, two–dimensional non-Abelian control over operation order built from the reduced Burau representation of the braid group $B_{3}$, specialised at $t=e^{i\omega}$ and unitarized by Squier’s Hermitian form. Coupled to two non-commuting qubit unitaries $A,B$, the resulting switch admits a closed expression for the single-shot Helstrom success and a fixed-order ceiling $p_{\rm fixed}^{*}$, yielding an explicit, analytic witness gap $\Delta(\omega)=p_{\rm switch}(\omega)-p_{\rm fixed}^{*}$. We prove that $\Delta(\omega)>0$ is achievable, thereby certifying causal non-separability by purely algebraic means, and confirm this behaviour numerically. Conceptually, this furnishes a minimal non-Abelian $B_{3}$ control for a Gedankenexperiment in anyonic statistics.
@@ 1 Introduction
Indefinite causal order (ICO) is typically modeled by the quantum switch- a coherent superposition of two operation orders, $AB$ and $BA$ [1, 2]. Most certifications of ICO employ the process-matrix framework and causal witnesses formulated as semidefinite programs (SDPs) [3, 4].
@@
Within this framework, the control mixer $M(\omega)$ is a $2\times 2$ unitary obtained from a braid word $w$, and the overall switch combines two non-commuting target operations $A,B\in U(2)$ through a controlled superposition of orders. The resulting process $T(\omega)$ admits closed-form discrimination probabilities through Helstrom’s theorem [13], allowing us to define an analytic witness gap $\Delta(\omega)$ that certifies causal non-separability when positive.
@@ 3 Braid group control and the switch
Let $B_{3}$ denote the braid group on $3$ strands. There are several equivalent definitions of this mathematical object.
@@ 5 Switch block and test device
$M(\omega)\ :=\ R(\omega)\,\beta(w)\big|_{s=e^{i\omega/2}}\,R^{-1}(\omega)\ \in U(2).$ (11)
@@
are U(2)U(2)–valued, real-analytic functions of $\omega$ on $I$.
@@
$J(\omega)$ has real-analytic entries in $\omega$, hence $\det J(\omega)$ is a real-analytic $2\pi$–periodic function; its zero set $Z$ is finite unless $\det J\equiv 0$ (which is not the case here). Thus $\Omega_{+}$ is a finite union of open intervals with endpoints in $Z$. On a compact subinterval of $I$, the spectrum of $J(\omega)$ is uniformly bounded away from 0, so the Cholesky factor $R(\omega)$ may be chosen to be real-analytic, cf. [22, Chapter II, §6.1–6.4]. Conjugation by such a matrix $R(\omega)$ preserves analyticity, and JJ–unitarity of $\beta_{3}(\cdot)$ implies $M_{\mathrm{pre/post}}(\omega)\in U(2)$. ∎
@@
By Lemma 5.1, $M_{\mathrm{pre/post}}(\omega)$ are real-analytic on $I$. The map $\omega\mapsto e^{i\theta(\omega)}$ is smooth and real-analytic, if $\theta$ is. Hence, so is $S(\theta(\omega))$. Products of real-analytic matrix-valued functions are real-analytic. ∎
@@ 6 Helstrom’s formula and convexity witness
$\Delta(\omega)\ :=\ p_{\rm switch}(\omega)-p_{\rm fixed}^{*}.$ (17)
@@ 7 Main Results
    the single-shot Helstrom success $p_{\rm switch}(\omega)$ for discriminating the switch process $S(\theta(\omega))$ from identity; and
@@
The witness gap $\Delta(\omega)=p_{\rm switch}(\omega)-p_{\rm fixed}^{*}$ for the test device $T(\omega)$ with the above parameters is strictly positive for $|\omega|<\frac{2\pi}{3}$.
@@
Figure 2: Left: verification of Squier’s $J$-unitarity for $\beta_{1}$, $\beta_{2}$. Center: Euclidean metric error of the unitarized word $U(\omega)$ within the positivity intervals of $J(\omega)$. Right: witness gap $\Delta(\omega)$ showing a positive region (non-convexity of causal order) where $J(\omega)\succ 0$.
Within the positive-definite window $|\omega|<2\pi/3$, the Squier generators satisfy $J$-unitarity to numerical precision better than $10^{-14}$, and the conjugated word $U(\omega)$ is Euclidean-unitary with error $\|U^{\dagger}U-I\|\!<\!10^{-14}$. The Helstrom success for the switch exceeds the fixed-order ceiling by a gap $\Delta_{\max}\approx 0.1796$ at $\omega\approx 1.754$, confirming causal non-separability.
@@
The control mixer $M(\omega)$ embodies the minimal non-Abelian structure of $B_{3}$: two non-commuting generators $\beta_{1}(\omega)$ and $\beta_{2}(\omega)$ acting unitarily on a two-dimensional control space. The existence of a positive witness gap $\Delta(\omega)>0$ demonstrates that the combined process cannot be expressed as a convex mixture of fixed-order operations.

### v2

@@ Abstract
We introduce a frequency–tunable, two–dimensional non-Abelian control of operation order constructed from the reduced Burau representation of the braid group $B_{3}$, specialised at $t=e^{i\omega}$ and unitarized by Squier’s Hermitian form. Coupled to two non-commuting qubit unitaries $A,B$, the resulting switch admits a closed expression for the single-shot Helstrom success probability and a fixed-order ceiling $p_{\mathrm{fixed}}$, defining the analytic witness gap $\Delta(\omega)=p_{\mathrm{switch}}(\omega)-p_{\mathrm{fixed}}$. The sign change of $\Delta(\omega)$ across the Squier positivity window reveals alternating constructive and destructive interference of causal orders, a hallmark of non-Abelian control, while $\Delta(\omega)>0$ certifies algebraic causal non-separability. Numerical simulations confirm both enhancement and suppression regimes, establishing a minimal $B_{3}$ braid control that reproduces the characteristic interference pattern expected from a Gedankenexperiment in anyonic statistics.
@@ 1 Introduction
Indefinite causal order (ICO) is typically modeled by the quantum switch- a coherent superposition of two operation orders, ABAB and BABA [1, 2]. Most certifications of ICO employ the process-matrix framework and causal witnesses formulated as semidefinite programs (SDPs) [3, 4].
@@
Within this framework, the control mixer $M(\omega)$ is a $2\times 2$ unitary obtained from a braid word $w$, and the overall switch combines two non-commuting target operations $A,B\in U(2)$ through a controlled superposition of orders. The resulting process $T(\omega)$ admits closed-form discrimination probabilities through Helstrom’s theorem [13], allowing us to define an analytic witness gap $\Delta(\omega)$ that certifies causal non-separability when positive, and makes distinction between the Abelian and non-Abelian cases by flipping its sign.
@@ 3 Braid group control and the switch
Let $B_{3}$ denote the braid group on 33 strands. There are several equivalent definitions of this mathematical object.
@@ 5 Switch block and test device
$M(\omega)\ :=\ R(\omega)\,\beta(w)\big|_{s=e^{i\omega/2}}\,R^{-1}(\omega)\ \in U(2),$ (11)
for some braid word $w\in B_{3}$.
@@
are U(2)U(2)–valued, real-analytic functions of $\omega$ on II.
@@
$J(\omega)$ has real-analytic entries in $\omega$, hence $\det J(\omega)$ is a real-analytic $2\pi$–periodic function; its zero set $Z$ is finite unless $\det J\equiv 0$ (which is not the case here). Thus $\Omega_{+}$ is a finite union of open intervals with endpoints in $Z$. On a compact subinterval of II, the spectrum of $J(\omega)$ is uniformly bounded away from 0, so the Cholesky factor $R(\omega)$ may be chosen to be real-analytic, cf. [22, Chapter II, §6.1–6.4]. Conjugation by such a matrix $R(\omega)$ preserves analyticity, and JJ–unitarity of $\beta_{3}(\cdot)$ implies $M_{\mathrm{pre/post}}(\omega)\in U(2)$. ∎
@@
By Lemma 5.1, $M_{\mathrm{pre/post}}(\omega)$ are real-analytic on II. The map $\omega\mapsto e^{i\theta(\omega)}$ is smooth and real-analytic, if $\theta$ is. Hence, so is $S(\theta(\omega))$. Products of real-analytic matrix-valued functions are real-analytic. ∎
@@ 6 Helstrom’s formula and convexity witness
and the full test device performance
$p_{\rm test}(\omega)\ :=\ p^{*}(I\otimes I,\ T(\theta(\omega))).$ (17)
@@
$\Delta(\omega)\ :=\ p_{\rm switch}(\omega)-p_{\rm fixed}^{*}$ (18)
for the switch $S(\omega)$ and, with a slight abuse of notation,
$\Delta(\omega)\ :=\ p_{\rm test}(\omega)-p_{\rm fixed}^{*}$ (19)
for the test device $T(\omega)$ complete with the mixers $M_{\rm pre/post}(\omega)$.
@@ 7 Main Results
    the single-shot Helstrom success probabilities $p_{\rm switch}(\omega)$ and $p_{\rm test}(\omega)$;
@@
The witness gap $\Delta(\omega)=p_{\rm switch}(\omega)-p_{\rm fixed}^{*}$ for the switch device $S(\omega)$ with the above parameters is strictly positive for $|\omega|<\frac{2\pi}{3}$. The witness gap for the test device $T(\omega)$ changes its sign.
@@
Figure 2: Left: Verification of Squier’s $J$-unitarity for $\beta_{1},\beta_{2}$.
Center: Euclidean metric error of $U(\omega)$ within the positivity window of $J(\omega)$.
Right: Witness gap $\Delta(\omega)$, both Abelian and non-Abelian, in the region where $J(\omega)\succ 0$.
Within the positive-definite window $|\omega|<2\pi/3$, the Squier generators satisfy $J$-unitarity to numerical precision better than $10^{-14}$, and the conjugated word $U(\omega)$ is Euclidean-unitary with error $\|U^{\dagger}U-I\|\!<\!10^{-14}$.
The Helstrom success for the switch exceeds the fixed-order ceiling by a gap $\Delta_{\max}\approx 0.1796$ at $\omega\approx 1.754$, confirming causal non-separability.
However, the non-Abelian test device produces a witness gap that changes its sign. Namely, we have $\Delta_{\min}\approx-0.2113$ at $\omega\approx 0.951$.
@@
The control mixer $M(\omega)\neq I$ embodies the minimal non-Abelian structure of $B_{3}$: two non-commuting generators $\beta_{1}(\omega)$ and $\beta_{2}(\omega)$ acting unitarily on a two-dimensional control space. The existence of a positive witness gap $\Delta(\omega)>0$ demonstrates that the combined process cannot be expressed as a convex mixture of fixed-order operations. The existence of a sign flip shows a clear distinction between the Abelian and non-Abelian cases.
@@
The non-Abelian $M(\omega)$ sometimes rotates the two order components into more overlapping subspaces, making them harder to distinguish, and thus the witness gap for the test device $T(\omega)$ changes it sign, while it stays always positive for the switch device $S(\omega)$.